OpenStudy (anonymous):

I am having a ton of difficulty seeing how to proceed when solving this partial fraction integration problem. Anyone have some tips/tricks when dealing with something like this??The integral of : (x^2-3x+7)/(x^2-4x+6)^2

OpenStudy (misty1212):

ick

OpenStudy (misty1212):

\[\frac{Ax+B}{x^2-4x+6}+\frac{Cx+D}{(x^2-4x+6)^2}\]

OpenStudy (anonymous):

First check if the polynomial in the denominator has real roots: \[\text{discriminant}=(-4)^2-4(1)(6)=16-24<0~~\implies~~\text{no real roots}\] which means the denominator is irreducible into linear terms. You'll have to settle for the form provided above.

OpenStudy (misty1212):

but it has been cooked up to give nice whole number solutions http://www.wolframalpha.com/input/?i=partial+fractions+%28x^2-3x%2B7%29%2F%28x^2-4x%2B6%29^2

OpenStudy (anonymous):

Yeah, I've seen the answer. But i'm trying to understand their logic and see how to proceed were I to get another beautiful question like this. Thanks for the help :)

OpenStudy (anonymous):

\[\begin{align*} \frac{x^2-3x+7}{(x^2-4x+6)^2}&=\frac{Ax+B}{x^2-4x+6}+\frac{Cx+D}{(x^2-4x+6)^2}\\\\ x^2-3x+7&=(Ax+B)(x^2-4x+6)+Cx+D\\\\ &=Ax^3+(-4A+B)x^2+(6A-4B+C)x+(6B+D) \end{align*}\] Matching up coefficients gives the system \[\begin{cases}A=0\\-4A+B=1\\6A-4B+C=-3\\6B+D=7\end{cases}\]

OpenStudy (anonymous):

Thanks a ton Sith! Thats beautiful. I've just completed that and i've integrated the first function \[1/(x^2-4x+6) \to \int\limits 1/((x-2)^2+2 ) \to (1/\sqrt{2})\arctan(x/\sqrt{2})\]

OpenStudy (anonymous):

currently working on \[\int\limits (x+1)/(x^2-4x+6)dx\]

OpenStudy (anonymous):

In your first integral, you're missing a small detail. \[\int\frac{du}{u^2+a^2}=\frac{1}{a}\arctan\frac{u}{a}\] but here \(u=x-2\), not \(x\).

OpenStudy (anonymous):

Thank you! Yes, and that is how I manage to lose tons of marks in exams....

OpenStudy (anonymous):

Np :) What are you trying for the second integral?

OpenStudy (anonymous):

I want to find a value for which the derivative is x+1 so I can cancel the numerator. Does that seem feasible?

OpenStudy (anonymous):

It is, but it's not the only way to approach it. In the end you'll have to use a trig sub anyway.

OpenStudy (anonymous):

Are there more efficient methods that you would recommend?

OpenStudy (anonymous):

Not for an indefinite integral, no. I'd go straight for a trig sub. Complete the square again, etc.

OpenStudy (anonymous):

Through manipulation of the numerator i think i've got it. If at all possible can you confirm this makes sense. \[\int\limits(x+1)/(x^2-4x+6)dx \to \int\limits (x-2+3)/(x^2-4x+6)dx\] through subsitution \[u=x^2-4x+6 \to du=2x-4dx \to 1/2du=x-2dx\] \[1/2\int\limits3du/u \to 3/2\int\limits du/u \to 3/2\ln \left| u \right| \to 3/2\ln \left| x^2-4x+6 \right|+C\]

OpenStudy (anonymous):

Keep in mind, the denominator is being squared!

OpenStudy (anonymous):

You just broke my heart. You're right again... I think I need a breather... making so many stupid mistakes. Alright I'll get back at it. You've been great. :)

OpenStudy (anonymous):

So just to confirm though, my manipulating of the numerator and everything I've done omitting the fact that the denom is squared was correct and not something "illegal".

OpenStudy (anonymous):

Right, your approach just uses the fact that \(a-a=0\) for all real \(a\). So \(b+a-a=b\) for all real \(b\).

OpenStudy (anonymous):

It seems there is a mistake nonetheless. I cannot assume that it will become 3du/u. The 3 should become its own integral since the three isn't a coefficient. As many mistakes as I am making, I am learning. :) Fall down, get back up again lol. Loads of falling down tonight.

OpenStudy (anonymous):

Thanks for the help SithsAndGiggles. I'll be moving on to the next question.

OpenStudy (anonymous):

Here's how I would have worked the squared-denominator integral: \[\int\frac{x+1}{(x^2-4x+6)^2}\,dx=\int\frac{x+1}{((x-2)^2+2)^2}\,dx\] Substitute \(\dfrac{x-2}{\sqrt2}=\tan t\), so \(x=\sqrt2\tan t+2\) and \(dx=\sec^2t\,dt\). \[\begin{align*}\int\frac{x+1}{((x-2)^2+2)^2}\,dx&=\int\frac{\sqrt2\tan t+3}{(2\tan^2t+2)^2}\sec^2t\,dt\\\\ &=\frac{1}{4}\int\frac{\sqrt2\tan t+3}{(\sec^2t)^2}\sec^2t\,dt\\\\ &=\frac{1}{4}\int\frac{\sqrt2\tan t+3}{\sec^2t}\,dt\\\\ &=\frac{\sqrt2}{4}\int \sin t\cos t\,dt+\frac{3}{4}\int \cos^2t\,dt\\\\ &=\frac{\sqrt2}{2}\int \sin2t\,dt+\frac{3}{8}\int (1+\cos2t)\,dt \end{align*}\]

OpenStudy (anonymous):

If I were to use your approach: \[\int\frac{x+1}{(x^2-4x+6)^2}\,dx=\frac{1}{2}\int\frac{2x-4}{(x^2-4x+6)^2}\,dx+\int\frac{3}{((x-2)^2+2)^2}\,dx\] Sub for the denominator in the first integral, and the previously used trig sub for the rightmost integral. \(y=x^2-4x+6\) so that \(dy=(2x-4)\,dx\); \(x-2=\sqrt2\tan z\), so \(dx=\sqrt 2\sec^2z\,dz\). \[\begin{align*}\int\frac{x+1}{(x^2-4x+6)^2}\,dx&=\frac{1}{2}\int\frac{dy}{y}+\frac{3}{4}\int\frac{\sec^2z}{(\tan^2z+1)^2}\,dz\\\\ &=\frac{1}{2}\int\frac{dy}{y}+\frac{3}{4}\int\cos^2z\,dz\\\\ &=\frac{1}{2}\int\frac{dy}{y}+\frac{3}{8}\int(1+\cos2z)\,dz \end{align*}\]

OpenStudy (anonymous):

Oops, missing a factor of \(\sqrt2\) in the second numerator, but you get the idea.

OpenStudy (anonymous):

Also, should be \(\dfrac{dy}{y^2}\) >_<

OpenStudy (anonymous):

Yeah. I'm sure they probably wanted us to use that considering we have just finished trig sub. The factor you're missing comes from your \[x=2√tant+2 \] you forgot the \[sqrt{(2)}\] in front of your sec. \[dx=sec2tdt.\]

OpenStudy (anonymous):

lol sorry about the format there....

OpenStudy (anonymous):

I find these questions hard due to the incredible amount of areas where you can make a silly mistake. :)

OpenStudy (anonymous):

THey are fun challenges though.

OpenStudy (anonymous):

Integrating rational functions via partial fractions can get pretty tedious, and it doesn't help that there's so much room to make mistakes...